A CF-18 jet fighter flying at 350m/s engages it’s afterburners and accelerate at a rate of 12.6 m/s² to a velocity of 600 m/s. How far does the fighter jet travel during acceleration?


1
Expert's answer
2021-10-06T08:15:16-0400

Find the time it takes to accelerate from 350 to 600 m/s:


"v=v_0+\alpha t,\\\space\\
t=\frac{v-v_0}{a}=19.84\text { s}."

The distance is


"d=v_0t+\frac{at^2}{2}=9424\text{ m}."